# TeSent: A Benchmark Dataset for Fairness-aware Explainable Sentiment Classification in Telugu

### v1

@@ 3. Data Collection & Preprocessing
To filter out non-Telugu content and code-mixed sentences from the collected corpus of 154,959 sentences, we utilized the UNICODE for Telugu and removed all sentences containing any non-Telugu fonts. However, we didn’t remove punctuation, emojis, etc., and the size of the final collected data was roughly 120,000. From this, we removed sentences that had only emojis or punctuation, and we were left with roughly 90,000. From this collected data, we removed duplicates in two steps.
@@
To address duplicates, we carried out a two-step deduplication process. First, we remove syntactically (near) duplicate sentences by computing the Jaccard similarity between bigrams of each sentence pair, retaining the first unique sentence, and discarding later ones that exceed a 75% similarity threshold value. This step reduced the dataset to 86,519 unique sentences. In the second step, we used mBERT embeddings to remove semantically (near)duplicate sentences. For each new sentence, it calculates the pairwise cosine similarity between its embedding and the embeddings of all previously accepted, non-duplicate sentences. If the maximum similarity among these comparisons is below a 90% threshold value, the sentence is considered sufficiently different and is kept, else discarded. After this semantic filtering, the dataset was reduced to 26,150 sentences. We next anonymized the dataset by replacing personally identifiable information such as email addresses, phone numbers, URLs, dates, etc, with applicable placeholders (<NUMBER>, <EMAIL>, <LINK>, <PHONE>, <PERCENT>, <DATE>, <TIME>, <MONEY>). After this step, we randomly selected 650 sentences for internal testing related to software validation and annotator recruitment, which left us with a working set of 25,500 sentences.
4. Annotation Process
4.1. Annotator Recruitment
[…]
4.2. Annotation Framework
In contrast to prior efforts, we adopt a distinct methodological framework for annotation. To support this, we also developed custom in-house annotation software tailored to our specific design principles, as detailed in section 4.2.1. Each annotator annotates every sentence following two layers of annotation. First, annotators are instructed to assign a primary sentiment label. Following this, they are asked to highlight specific words or phrases in the sentence that serve as human rationales for their chosen label. To capture subtler sentiment dynamics, annotators are then given the option to provide a secondary sentiment label if they believe one reasonably coexists with the primary annotation. However, for brevity and to reduce cognitive load, we do not collect rationales for this secondary label.
4.2.1. Software Design
@@
The custom-built annotation interface included several features tailored to our design principles. Annotators logged in securely using unique credentials and were presented with one sentence at a time. For each sentence, they were required to select a primary sentiment label: Positive, Negative, or Neutral, and had the option to indicate an alternative sentiment if applicable. A validation check is flagged in case of identical primary and alternative labels. If any annotator finds a sentence to be grammatically incorrect, incomplete, off-topic, or uninterpretable, s/he will have an option to flag the same as “Mark as Bad Sentences” on the fly. At least one such flag would mark the sentence as Invalid and exclude it from the dataset.
@@
On the interface, rationale selection is done by double-clicking the word(s) from the sentence. These rationales, along with annotations, were stored in a MongoDB database. To ensure consistency and minimize fatigue, the dataset was split into 17 pools of 10 batches each (150 sentences per batch). The annotation workload was distributed systematically to ensure reliability and reduce annotator fatigue. A progress bar, gamified leaderboard, and periodic manual reminders were used to maintain annotator engagement and ensure timely batch completion. A snapshot of the annotation interface is attached in Appendix E.
From the annotated corpus of 25,500 sentences, 23,144 were found to be valid. In addition, out of 650 sentences selected for testing, 517 were valid. This brought the total number of valid annotated sentences to 23,661.
4.3. Finalizing the Label
[…]
Out of 23,661 total valid and annotated sentences, 21,018 had only primary annotations from all annotators, of which 20,313 had a clear majority winner and were retained. The remaining 2,643 sentences included at least one annotation with both a primary and secondary label; after removing 61 due to a technical glitch, 2,582 remained. From these, annotations with ¡+, -¿ pairs were removed where present, and majority voting over the priority labels yielded 2,272 sentences. This resulted in a corpus of 22,585 sentences. Finally, Telugu-speaking authors manually inspected the entire annotated dataset and, at their discretion, removed an additional 80 malformed or duplicate sentences, leading to a final total of 22,505. Overall, we retained over 96.5% of high-quality sentences with the final label from the set of valid sentences.
@@ 4.3. Finalizing the Label
Figure 1. Word Count Frequency for TeSent
@@
Figure 2. Category Wordcloud for TeSent
4.4. Annotation Statistics and Interpretation
Our final dataset consists of 44.99% neutral, 26.64% positive, and 28.37% negative instances. It has been annotated by a large group with varying levels of contribution. For an overview, we refer readers to Figure 4.
To assess label consistency among annotators, we measured inter-rater reliability using Krippendorff’s Alpha ($\alpha$) and Fleiss’s Kappa ($\kappa$) (Gwet, 2014). When considering all three classes, the agreement scores were $\alpha=0.4797$ and $\kappa=0.4796$, indicating moderate agreement.
However, we acknowledge that the Neutral class may introduce semantic ambiguity, especially when sentences exhibit overlapping sentiment cues. To account for this, we recalculated agreement after excluding Neutral-labelled sentences, which resulted in a substantially higher $\alpha=0.8992$.
This suggests that annotator agreement is much stronger when distinguishing strictly between Positive and Negative sentiment. To further illustrate the distribution of annotator agreement levels, we refer readers to Figure 3, which presents the cumulative distribution of how often one annotator agrees with the other two. As shown,  80% of annotators exhibit agreement levels above 82%, indicating a high degree of consistency across most of the dataset. However, a certain portion of the data remains ambiguous, likely due to the inherently fuzzy boundaries between sentiment categories, particularly in distinguishing Neutral sentiment from positive or negative classes.
Figure 3. Distribution of Agreement Levels
Figure 4. Annotation Distribution
5. Fairness Benchmarking
Sentiment classification systems can perpetuate various societal biases, including gender, racial, and identity-based biases. The presence of such bias often leads to disparities in classification accuracy or sentiment polarity across different demographic groups. Kiritchenko et al. (Kiritchenko and Mohammad, 2018) proposed the Equity Evaluation Corpus (EEC), an English dataset designed to evaluate age and racial bias across multiple NLP tasks related to sentiment and emotion. We extend and tune EEC accommodating niche morphological nuances in Telugu language to create TeEEC, Equity Evaluation Corpus in Telugu, and propose a robust fairness evaluation framework for sentiment classification in Telugu, with a focus on gender and religion bias. Although our bias evaluation framework is specifically designed for sentiment classification, TeEEC can be applied to any NLP task involving sentiment or emotion.
5.1. TeEEC
Kiritchenko et al. (Kiritchenko and Mohammad, 2018) provided 11 English sentence templates categorized into two types: (a) with emotion words and (b) without emotion words. Examples include:
Type (a): ¡Person¿ feels ¡emotion word¿.
Type (b): I saw ¡person¿ in the market.
Type (a) consists of seven and Type (b) consists of four templates. Common emotion words were chosen from each of the following themes: Anger, Fear, Joy and Sadness. Emotion words of the first three categories exhibit strong negative emotion and the rest exhibit strong positive emotion. ¡emotion word¿s in the type (a) templates are replaced by those words, which decide the sentiment polarity of the sentences. ¡Person¿ and ¡person¿ placeholders can be replaced by name or noun phrase, which is indicative of the demographic group that (s)he belongs. For an example, the girl and the boy can represent a male and a female respectively in context of evaluating gender bias. For brevity, we denote both of ¡Person¿ and ¡person¿ as ¡person¿.
[…]
5.2. Evaluation Metrics
The evaluation matrices we employ aim to quantify bias in two distinct ways: (i) by measuring differences in predicted sentiment polarities across demographic groups using counterfactual test (Kiritchenko and Mohammad, 2018; Goldfarb-tarrant et al., 2023), and (ii) by assessing fairness in terms of equalized odds (Hardt et al., 2016), which is based on the difference in classification performance across groups.
5.2.1. Counterfactual Test
[…]
@@ 5.2.2. Equalized Odds (EO)
6. Explainability Benchmarking
Commonly used post-hoc local explainers can be broadly categorized into two types: perturbation-based and gradient-based explainers (Arrieta et al., 2020). In this work, we consider two widely adopted perturbation-based, model-agnostic explainers: LIME (Ribeiro et al., 2016) and SHAP (Lundberg and Lee, 2017), using the default selection of partition shap for SHAP.
Among the gradient-based methods, we include Gradient (Grad.) (Simonyan et al., 2013), Integrated Gradients (Int. Grad.) (Sundararajan et al., 2017), and their input-scaled variants: Gradient × Input (Grad. × Input) and Integrated Gradients × Input (Int. Grad. × Input).
To evaluate faithfulness, we compare our results using well-established erasure-based metrics: comprehensiveness (C($\uparrow$)), sufficiency (S($\downarrow$)) (DeYoung et al., 2020), and correlation with Leave-One-Out (L($\uparrow$)) scores (Jain and Wallace, 2019).
For plausibility analysis, we followed DeYoung et al. (DeYoung et al., 2020) and report token-level Intersection-Over-Union (I($\uparrow$)), token-level F1 scores (F($\uparrow$)), and Area Under the Precision-Recall Curve (A($\uparrow$)). For all measures, $\uparrow$’ signifies the more the better and $\downarrow$’ signifies the lower the better.
We conduct all experiments using the ferret-xai library with its default (hyper)parameter settings (Attanasio et al., 2023) and report the mean values for each measure.
7. Experimental Results
7.1. Model Training
We evaluate five state-of-the-art transformer models that support Telugu: mBERT (Devlin et al., 2019), XLM-R (Conneau et al., 2019), MuRIL (Khanuja et al., 2021), IndicBERT (Joshi, 2022) and TeBERT (Joshi, 2022). Detailed description of the models can be found in Appendix C. Each model is trained in two settings: first, without rationales(denoted as wo/r), and then with rationales (denoted as w/r). Training a model without rationales allows it to learn any arbitrary class of functions, as it simply tries to fit the training data without any guidance on how its decisions align with human reasoning. In contrast, training with rationales explicitly curates the learning process by incorporating human-provided reasoning into the training. This allows the model to better align with human values during training. Previous studies (Mathew et al., 2021) have demonstrated that this approach helps improve the fairness of the model by providing additional supervisory signals that guide the model’s decision-making process and reduce biases in its outputs. For the same, we adopt a specialized loss function, following the approach proposed by Zhang et al. (Zhong et al., 2019) and Pruthi et al. (Pruthi et al., 2022): we took the mean of the rationale vector, added a small constant ($1\mathrm{e}{-8}$) in the place of zeroes (0), and converted it into a single probability distribution ($\hat{A}$) using softmax. Next, we computed the attention from the [CLS] token to other tokens in the last layer, averaged across all attention heads ($A$). The new modified loss function for this training is [^4] For neutral sentences with no rationale, we only compute the CE loss.:
$\text{Modified\ Loss}=\text{CE Loss}+\lambda\cdot\text{KL}(A\parallel\hat{A})$
We refer readers to Table 1 for a consolidated overview of the models’ performance. For a detailed overview of the selection of hyperparameters for both types of training, we refer readers to Appendix D.
Model Accuracy F1 AUROC
mBERT wo/r 0.647 0.647 0.822
w/r 0.673 0.668 0.839
IndicBERT wo/r 0.737 0.736 0.883
w/r 0.734 0.735 0.893
MuRIL wo/r 0.710 0.708 0.870
w/r 0.722 0.722 0.885
XLM-R wo/r 0.711 0.711 0.865
w/r 0.717 0.716 0.877
Te-BERT wo/r 0.736 0.736 0.877
w/r 0.734 0.735 0.890
Table 1. Models’ performance for both types of training. wo/r: without rational, w/r: with rational
@@ 7.1. Model Training
Plausibility
mBERT wo/r 0.583 0.410 0.290 0.572 0.406 0.282 0.495 0.337 0.228 0.488 0.332 0.224 0.503 0.348 0.235 0.553 0.391 0.271
w/r 0.628 0.453 0.325 0.630 0.455 0.323 0.529 0.369 0.250 0.520 0.362 0.245 0.533 0.372 0.251 0.585 0.423 0.294
IndicBERT wo/r 0.626 0.464 0.333 0.570 0.416 0.289 0.510 0.391 0.276 0.517 0.396 0.280 0.520 0.388 0.273 0.570 0.421 0.309
w/r 0.635 0.479 0.338 0.598 0.437 0.315 0.520 0.398 0.282 0.531 0.410 0.287 0.540 0.415 0.290 0.585 0.440 0.318
MuRIL wo/r 0.609 0.457 0.326 0.591 0.444 0.304 0.512 0.387 0.267 0.505 0.378 0.260 0.524 0.391 0.268 0.578 0.419 0.296
w/r 0.622 0.469 0.338 0.605 0.453 0.318 0.518 0.394 0.277 0.510 0.378 0.263 0.522 0.387 0.269 0.580 0.423 0.301
XLM-R wo/r 0.572 0.399 0.274 0.560 0.391 0.269 0.495 0.341 0.227 0.490 0.335 0.219 0.510 0.348 0.229 0.570 0.385 0.263
w/r 0.609 0.433 0.305 0.585 0.414 0.289 0.515 0.378 0.254 0.505 0.368 0.247 0.518 0.380 0.253 0.574 0.397 0.277
Te-BERT wo/r 0.603 0.454 0.324 0.577 0.433 0.307 0.505 0.384 0.266 0.507 0.375 0.260 0.520 0.388 0.267 0.570 0.413 0.291
w/r 0.612 0.464 0.336 0.591 0.451 0.323 0.516 0.388 0.276 0.507 0.373 0.263 0.523 0.382 0.269 0.576 0.420 0.298
@@
Faithfulness
mBERT wo/r 0.457 -0.003 0.268 0.481 -0.011 0.296 0.305 0.126 0.077 0.125 0.330 -0.042 0.203 0.269 0.023 0.336 0.137 0.116
w/r 0.435 0.018 0.257 0.464 0.012 0.293 0.277 0.131 0.062 0.121 0.311 -0.040 0.145 0.295 -0.028 0.341 0.118 0.129
IndicBERT wo/r 0.474 0.059 0.273 0.529 0.029 0.419 0.370 0.116 0.043 0.210 0.295 0.023 0.192 0.319 -0.012 0.428 0.102 0.179
w/r 0.434 0.076 0.281 0.483 0.049 0.432 0.362 0.166 0.053 0.223 0.274 0.025 0.173 0.337 -0.021 0.396 0.110 0.164
MuRIL wo/r 0.474 0.021 0.284 0.481 0.029 0.308 0.345 0.115 0.056 0.114 0.352 -0.062 0.120 0.360 -0.008 0.441 0.048 0.227
w/r 0.458 0.023 0.289 0.465 0.026 0.294 0.338 0.106 0.045 0.135 0.305 -0.050 0.151 0.273 -0.008 0.430 0.031 0.265
XLM-R wo/r 0.468 -0.013 0.313 0.494 -0.017 0.383 0.321 0.098 0.082 0.078 0.347 -0.111 0.160 0.248 0.027 0.362 0.058 0.154
w/r 0.457 -0.013 0.304 0.485 -0.018 0.374 0.310 0.094 0.085 0.082 0.323 -0.124 0.126 0.272 0.003 0.366 0.045 0.173
Te-BERT wo/r 0.440 0.049 0.254 0.473 0.030 0.314 0.324 0.117 0.059 0.095 0.346 -0.082 0.141 0.283 0.000 0.418 0.042 0.249
w/r 0.454 0.060 0.265 0.480 0.043 0.315 0.350 0.113 0.056 0.120 0.358 -0.055 0.177 0.293 0.001 0.427 0.050 0.221
@@
Table 2. Combined Plausibility and Faithfulness Results. wo/r: without rational, w/r: with rational
7.2. Fairness Results
Table 3 and Table 4 present the results of the fairness evaluation with respect to gender and religion, respectively. In the context of gender, males are considered the privileged group; similarly, for religion, Hindus are regarded as the privileged group based on the demographics of native Telugu speakers [^5] https://en.wikipedia.org/wiki/Telugu_people/. So, we consider fairness evaluation between Hindu–Muslim and Hindu–Christian pairs. Negative $PDS$ values observed in three models in Table 3 indicate that the underprivileged gender, namely female, exhibits higher positive polarity compared to the privileged gender, male, within these models. In case of religion, the near absence of negative $PDS$ values indicates that the privileged religious group consistently exhibits higher positive polarity. Although training with rationales does not show consistent improvements in $PDS$ for gender—except in the case of XLM-R—significant decrease in $PDS$ are observed for both religious demographic pairs. For mBERT, IndicBERT, and Te-BERT, both $TPR_{diff}$ and $FPR_{diff}$ decrease in the gender bias evaluation when trained with rationale, whereas the differences are negligible for the remaining two models. For mBERT and IndicBERT, both $TPR_{diff}$ and $FPR_{diff}$ decrease significantly across both religious demographic pairs when trained with rationale, whereas the differences remain insignificant for the other two models. Mostly, in those few cases where the biases increase when trained with rationales, the increase is insignificant. Overall, our results indicate that training with rationale may reduce biases related to both gender and religion. Our results are consistent with the findings of Mathew et al. (Mathew et al., 2021), who reported similar bias patterns in hate speech classification for English. To fully understand the biases in our Telugu sentiment classification models, it is important to examine the biases in the embedding space of the pre-trained models, which is beyond the scope of this study. Assessing bias within the embedding space of pre-trained models for Telugu has received relatively little attention in NLP community. Since Telugu is an extremely low-resource language, the pre-trained models have been trained on much less data than those for high-resource languages. As a result, the scale and diversity of the biases in Telugu models may differ from what we see in English and other high-resource languages.
Model $PDS$ $TPR_{diff}$ $FPR_{diff}$
mBERT wo/r 0.004 0.008 0.004
w/r 0.008 0.003 0.002
IndicBERT wo/r 0.013 0.014 0.007
w/r 0.001 0.001 0.000
MuRIL wo/r -0.002 0.004 0.002
w/r 0.013 0.005 0.002
XLM-R wo/r -0.004 0.001 0.000
w/r -0.005 0.004 0.002
Te-BERT wo/r 0.000 0.010 0.005
w/r 0.003 0.001 0.001
Table 3. Gender bias evaluation results. wo/r: without rational, w/r: with rational
Model Hindi-Christian Hindu-Muslim
$PDS$ $TPR_{diff}$ $FPR_{diff}$ $PDS$ $TPR_{diff}$ $FPR_{diff}$
mBERT wo/r 0.067 0.046 0.023 0.092 0.048 0.232
w/r 0.024 0.011 0.005 0.053 0.016 0.008
IndicBERT wo/r 0.030 0.021 0.010 0.030 0.026 0.013
w/r 0.006 0.011 0.072 0.014 0.003 0.002
MuRIL wo/r 0.023 0.004 0.002 0.023 0.002 0.001
w/r 0.009 0.001 0.000 0.001 0.003 0.001
XLM-R wo/r 0.000 0.001 0.001 -0.006 0.005 0.002
w/r 0.006 0.004 0.002 0.007 0.005 0.002
Te-BERT wo/r 0.014 0.004 0.002 0.022 0.005 0.003
w/r 0.011 0.005 0.003 0.013 0.005 0.002
Table 4. Religion bias evaluation results. wo/r: without rational, w/r: with rational
7.3. Plausibility and Faithfulness Analysis Results
Table 2 contains plausibility and faithfulness analysis results for all models, respectively. Roughly, plausibility is the extent to which an explanation aligns with human judgment or expectations. From Table 2, it is evident that all types of models, trained with rationale supervision, substantially improved plausibility across all considered metrics. Furthermore, among the explainers, SHAP and LIME generally show stronger alignment with human rationales compared to gradient-based methods, though Int Grad × Input consistently holds up as a robust explainer across models.
However, unlike plausibility, we do not find any noticeable pattern in faithfulness analysis, as evident in Table 2. It seems that rationale supervision slightly alters the faithfulness behaviour, but the effect varies by model and attribution technique. For instance, SHAP and LIME, for the IndicBERT model, generally exhibit higher C (↑) and L (↑) values across. On the other hand, gradient-based methods—especially Grad × Input—tend to produce more negative or lower L (↑) scores in all cases other than IndicBERT, indicating less consistent faithfulness. Faithfulness is loosely defined as how well an explainer reflects the underlying reasoning of a model, and training with rationales is a specialised training mechanism meant to make the model align with human reasoning. However, since the explainers are essentially non-trainable objects and not the model itself (Rudin, 2019), these two have no causal relationship, and we believe any pattern, even if it exists, should be perpetually coincidental—which, however, is not the case in our study.
8. Conclusion and Future Direction
In this work, we introduced TeSent, a comprehensive benchmark dataset for Telugu sentiment classification, consisting of 22,505 sentences with multiple human-annotated rationales. TeSent supports extensive evaluation of explainability and fairness in sentiment classification. Through systematic experiments, we showed that training models with rationales improves classification accuracy, reduces model bias, and enhances alignment with human reasoning. We also curated TeEEC, a dedicated Equity Evaluation Corpus for Telugu. Using TeEEC, we performed detailed bias evaluations, alongside exhaustive evaluation suites for both plausibility and faithfulness of model explanations. All resources—including the TeSent and TeEEC datasets, as well as rationale-based and baseline models—are publicly released to facilitate further research in this under-represented space.
For future work, we aim to mechanistically interpret the models to better understand how training with rationales influences what the models learn—specifically, which underlying concepts are learned differently. We are particularly interested in how these concepts are represented, stored, and shared within multilingual models, especially in relation to dominant languages like English. Additionally, we plan to extend our bias evaluation by decomposing the shared representation space, enabling a more nuanced analysis of how bias manifests across languages.
Acknowledgements.
We are thankful to all the annotators, without whose effort and feedback this work would not have been possible. We also thank SRM University AP for funding the project. This work was supported by the UG-PG Seed Grant Scheme of SRM University AP, India, under Grant No. SRMAP/UG-PG/SEED/2024-25/30.
Appendix A Data Split
Figure 5 illustrates the distribution of data sources used in our dataset, with the majority coming from YouTube Comments (YT) (51.2%), followed by blogs comments (Blogs), news headlines (News), and Facebook & Instagram comments (FB-IG).
Figure 5. Data Split
Appendix B Sampling Algorithm for Generating Counterfactuals
B.1. Gender Bias

### v2

@@ Appendix C Preprocessing Pipeline
This section details the preprocessing steps applied to the raw scraped corpus prior to annotation. To filter out non-Telugu content and code-mixed sentences from the collected corpus of 154,959 sentences, we utilized the UNICODE for Telugu and removed all sentences containing any non-Telugu fonts. However, we didn’t remove punctuation, emojis, etc., and the size of the final collected data was roughly 120,000. From this, we removed sentences that had only emojis or punctuation, and we were left with roughly 90,000. From this collected data, we removed duplicates in two steps.
@@
To address duplicates, we carried out a two-step deduplication process. First, we remove syntactically (near) duplicate sentences by computing the Jaccard similarity between bigrams of each sentence pair, retaining the first unique sentence, and discarding later ones that exceed a 75% similarity threshold value. This step reduced the dataset to 86,519 unique sentences. In the second step, we used mBERT embeddings to remove semantically (near)duplicate sentences. For each new sentence, it calculates the pairwise cosine similarity between its embedding and the embeddings of all previously accepted, non-duplicate sentences. If the maximum similarity among these comparisons is below a 90% threshold value, the sentence is considered sufficiently different and is kept, else discarded. After this semantic filtering, the dataset was reduced to 24,593 sentences. We next anonymized the dataset by replacing personally identifiable information such as email addresses, phone numbers, URLs, dates, etc, with applicable placeholders (<NUMBER>, <EMAIL>, <LINK>, <PHONE>, <PERCENT>, <DATE>, <TIME>, <MONEY>). After this step, we randomly selected 607 sentences for internal testing related to software validation and annotator recruitment, which left us with a working set of 23,986 sentences.
Appendix D Annotator Recruitment
This section describes the recruitment, screening, and compensation of annotators involved in the sentiment and rationale annotation task. First, we issued an open call through internal mailing lists, social media, and the internal organization networks of the organization, targeting native Telugu speakers. Prospective annotators (who are native speakers) were asked to complete a linguistic proficiency survey via Google Form. Participants labelled the sentiment and the rationales of a few selected Telugu sentences as Positive, Negative, or Neutral. Submissions were manually reviewed for accuracy, consistency, and quality of rationale selection by the Telugu-speaking authors against each sentiment. Out of 112 responses, based on the overall performance, a total of 95 annotators were selected by the Telugu-speaking authors for the main task. Next, a pilot annotation task was conducted among the Telugu-speaking authors to finalize the interface for our annotation software for web browsers for both mobile and desktop versions (such as whether “double-click" or “single-click" for selecting rationale would be more effective). We remunerated all annotators for completing each batch of 150 sentences. However, among selected annotators, 21 defaulted without completing a single batch by the end of the annotation phase, leaving us a total of 74 active annotators.
Appendix E Annotation Software Design
@@
This section details the design principles and features of the custom-built annotation interface used in this study. The custom-built annotation interface included several features tailored to our design principles. Annotators logged in securely using unique credentials and were presented with one sentence at a time. For each sentence, they were required to select a primary sentiment label: Positive, Negative, or Neutral, and had the option to indicate an alternative sentiment if applicable. A validation check is flagged in case of identical primary and alternative labels. If any annotator finds a sentence to be grammatically incorrect, incomplete, off-topic, or uninterpretable, s/he will have an option to flag the same as “Mark as Bad Sentences” on the fly. At least one such flag would mark the sentence as Invalid and exclude it from the dataset.
@@
On the interface, rationale selection is done by double-clicking the word(s) from the sentence. These rationales, along with annotations, were stored in a MongoDB database. To ensure consistency and minimize fatigue, the dataset was split into 17 pools of 10 batches each (150 sentences per batch). The annotation workload was distributed systematically to ensure reliability and reduce annotator fatigue. A progress bar, gamified leaderboard, and periodic manual reminders were used to maintain annotator engagement and ensure timely batch completion. A snapshot of the annotation interface is attached in Appendix L.
Appendix F Distribution of Word-count, Genre and Annotator Contributions
@@
Figure 3: Word Count Frequency for TeSent
@@
Figure 4: Category Wordcloud for TeSent
Distribution of annotation contributions across annotators in the TeSent dataset, showing the number of instances annotated per annotator is given in Figure 5.
Figure 5: Annotation Distribution
Appendix G Descriptions of the Fairness Metrics
G.1 Counterfactual Test
Counterfactual sentence pairs consist of two sentences, each drawn from the respective sets corresponding to the two demographic groups. Each counterfactual sentence represents one of the three sentiment classes $\in\text{\{negative, positive, neutral\}}$. Counterfactual pairs for the neutral class are generated using type (b) sentences by replacing the <person> placeholder with names and noun phrases taken from the name corpora of the two demographic groups. Since there is no one-to-one correspondence between names across demographic groups, counterfactual pairs are formed by considering all possible combinations of names, with one name drawn from each group. The 10 noun phrases exhibit a one-to-one correspondence across demographic groups, resulting in 10 counterfactual pairs for each type (b) template. Counterfactual pairs for the negative and positive class are generated using the type (a) sentences. For both the positive and negative classes, counterfactual pairs are constructed by replacing the <person> placeholder with names and noun phrases from the two demographic groups, for each <emotion word> corresponding to the respective sentiment polarity, as previously described. Given the large number of possible counterfactual pairs, we employ a randomized sampling algorithm to select 1320 counterfactuals for the case of gender, and 1200 counterfactuals for the case of religion, balanced across the three sentiment classes. The reported results are averaged over ten runs using different random seeds. The sampling algorithm is detailed in Appendix I. The counterfactual test measures the Polarity Difference Score (PDS), described as follows.
@@ G.2 Equalized Odds (EO)
Appendix H Detailed Result
Table 2 and Table 3 provide the results for gender and religion bias evaluation respectively. Table 4 provides Plausibility and Faithfulness evaluation Results.
Model PDSPDS $TPR_{diff}$ $FPR_{diff}$
mBERT wo/r -0.004 0.022 0.010
w/r -0.004 0.025 0.013
IndicBERT wo/r 0.003 0.006 0.003
w/r 0.003 0.003 0.003
MuRIL wo/r 0.008 0.011 0.006
w/r 0.008 0.010 0.004
XLM-R wo/r -0.004 0.017 0.007
w/r -0.007 0.008 0.004
Te-BERT wo/r 0.005 0.005 0.004
w/r -0.002 0.005 0.004
Table 2: Gender bias evaluation results. wo/r: without rational, w/r: with rational
Model Hindu-Christian Hindu-Muslim
PDSPDS $TPR_{diff}$ $FPR_{diff}$ PDSPDS $TPR_{diff}$ $FPR_{diff}$
mBERT wo/r 0.023 0.012 0.014 0.045 0.025 0.029
w/r 0.035 0.022 0.016 0.063 0.034 0.033
IndicBERT wo/r 0.004 0.019 0.013 0.012 0.011 0.004
w/r 0.020 0.013 0.008 0.012 0.007 0.004
MuRIL wo/r 0.048 0.033 0.020 0.051 0.036 0.020
w/r 0.017 0.012 0.006 0.014 0.013 0.006
XLM-R wo/r 0.004 0.015 0.007 0.012 0.010 0.006
w/r 0.014 0.012 0.005 0.012 0.011 0.005
Te-BERT wo/r 0.003 0.005 0.004 0.010 0.006 0.006
w/r 0.002 0.006 0.003 0.020 0.013 0.006
Table 3: Religion bias evaluation results. wo/r: without rational, w/r: with rational
@@ Appendix H Detailed Result
Plausibility mBERT wo/r 0.568 0.396 0.277 0.565 0.392 0.271 0.549 0.370 0.257 0.409 0.262 0.167 0.422 0.281 0.180 0.522 0.363 0.247
w/r 0.570 0.396 0.278 0.572 0.405 0.281 0.556 0.386 0.270 0.403 0.257 0.164 0.425 0.277 0.176 0.507 0.348 0.232
IndicBERT wo/r 0.555 0.419 0.294 0.549 0.403 0.277 0.548 0.444 0.317 0.392 0.308 0.208 0.410 0.315 0.214 0.572 0.433 0.308
w/r 0.576 0.430 0.305 0.562 0.414 0.288 0.583 0.463 0.334 0.416 0.329 0.225 0.442 0.329 0.225 0.573 0.463 0.298
MuRIL wo/r 0.585 0.446 0.319 0.594 0.447 0.318 0.566 0.436 0.31 0.424 0.309 0.207 0.421 0.295 0.197 0.584 0.435 0.31
w/r 0.598 0.448 0.32 0.607 0.451 0.321 0.567 0.447 0.32 0.415 0.303 0.202 0.425 0.311 0.209 0.6 0.449 0.32
XLM-R wo/r 0.577 0.43 0.303 0.588 0.438 0.307 0.56 0.424 0.298 0.407 0.296 0.195 0.395 0.271 0.177 0.508 0.378 0.259
w/r 0.575 0.429 0.303 0.582 0.432 0.303 0.554 0.427 0.303 0.38 0.275 0.181 0.395 0.291 0.191 0.477 0.367 0.249
Te-BERT wo/r 0.578 0.437 0.309 0.595 0.448 0.316 0.551 0.422 0.298 0.385 0.277 0.187 0.418 0.281 0.186 0.596 0.445 0.318
w/r 0.583 0.437 0.31 0.604 0.459 0.327 0.557 0.423 0.299 0.386 0.271 0.182 0.448 0.323 0.217 0.614 0.454 0.329
@@
Faithfulness mBERT wo/r 0.426 -0.021 0.239 0.456 -0.036 0.281 0.267 0.101 0.062 0.118 0.275 -0.033 0.147 0.262 -0.01 0.35 0.059 0.156
w/r 0.442 0.008 0.232 0.479 -0.019 0.273 0.3 0.12 0.064 0.121 0.329 -0.05 0.164 0.304 -0.004 0.342 0.136 0.123
IndicBERT wo/r 0.428 0.061 0.212 0.505 0.032 0.359 0.346 0.16 0.038 0.209 0.281 0.019 0.188 0.301 0.001 0.397 0.101 0.146
w/r 0.479 0.05 0.261 0.562 0.029 0.413 0.376 0.158 0.055 0.217 0.282 0.039 0.199 0.306 0.025 0.437 0.108 0.168
MuRIL wo/r 0.486 0.036 0.265 0.505 0.03 0.307 0.352 0.119 0.053 0.117 0.352 -0.073 0.156 0.316 -0.016 0.461 0.038 0.217
w/r 0.454 0.011 0.253 0.473 0.007 0.274 0.334 0.117 0.044 0.094 0.34 -0.077 0.125 0.311 -0.015 0.439 0.02 0.235
XLM-R wo/r 0.481 -0.014 0.271 0.512 -0.025 0.35 0.332 0.097 0.075 0.095 0.345 -0.115 0.119 0.294 -0.02 0.364 0.048 0.155
w/r 0.481 -0.027 0.299 0.5 -0.029 0.374 0.309 0.112 0.073 0.07 0.359 -0.121 0.135 0.267 0.01 0.331 0.09 0.133
Te-BERT wo/r 0.455 0.046 0.26 0.483 0.035 0.319 0.326 0.123 0.04 0.096 0.341 -0.083 0.113 0.356 -0.035 0.439 0.041 0.237
w/r 0.451 0.026 0.265 0.476 0.017 0.319 0.322 0.118 0.024 0.097 0.346 -0.062 0.161 0.267 0.021 0.417 0.034 0.212
@@
Table 4: Combined Plausibility and Faithfulness Results. wo/r: without rational, w/r: with rational
Appendix I Sampling Algorithm for Generating Counterfactuals
I.1 Gender Bias